# ScrollTimes: Tracing the Provenance of Paintings as a Window into History

### v1

@@ 6.2 Art Historian Reviews
We evaluated the efficacy and usability of ScrollTimes by interviewing experts ( Section 3) online, who were not co-authors of this manuscript.
Procedure. Each 90-minute interview consisted of a 5-minute consent and demographic questionnaire, a 20-minute introduction to our motivation and system design, a 45-minute free exploration of ScrollTimes, and a 20-minute evaluation of its effectiveness and usability.
Effectiveness of the system. All experts agreed that ScrollTimes would enhance their research. A1-A5 found it particularly useful for improving efficiency in seal identification and inscription analysis. A4 mentioned their current workflow was still in the “manual workshop” stage, involving searching different platforms to confirm the seal owner, leading to a time-consuming and inefficient process. This tedious and repetitive work adversely affects their research efficiency and enthusiasm. The automatic method provided by ScrollTimes can alleviate the laborious seal extraction process, making it a productivity tool that allows them to focus on other analytical tasks. A3 found that the combination of multiple databases saves time searching for background information on figures in paintings, and ScrollTimes can help them quickly find research entry points. A4 and A5 suggested using ScrollTimes as an auxiliary tool in the classroom of art history, improving teaching efficiency and student learning interest. E1 and E2 found the social relationship information reflected in the inscriptions helpful. E1 suggested the inscriptions could be a reliable data source to supplement the CBDB database, verifying existing data and expanding the data source, breaking down barriers between painting and text data.
Usability of the system. All experts agreed that the view designs are reasonable, and the exploration process is clear. A3-5 found the circular design of the handscroll (Fig. 3E1) to be “eye-catching.” Although the circular design may compromise the original aesthetic beauty of the handscroll, A4 believed that it allows viewers to see the whole picture of the handscroll at a glance, and the interaction with the circular design does not lose any original information. A3 and A5 approved of the word cloud design in the circular design, believing it provides a good summary of seal and inscription information and can inspire analysis. The comparative presentation in the Seals and Inscriptions View (Fig. 3E1-3) enhances the results’ trustworthiness. The Linked Data View (Fig. 3E2) provides rich person-to-person association information that the experts welcomed. While the Cohort View (Fig. 3E2-2) was familiar to E1-2, art historians expressed the need for some learning costs. A1 and A4 found the Biography View to be the most impressive, stating that “using a familiar scroll as a metaphor for life is excellent.” The coding on the life-scroll is rich but not difficult to understand.
@@ 7 Discussion
This section summarizes the lessons learned from working with art historians (A1-A5) and CBDB editors (E1 and E2), as well as the limitations and future work of ScrollTimes.
@@ 7.1 Design Implications
Open-mindedness. In addition to being a carrier of historical information, the handscroll is also a traditional work of art. Out of respect for ancient artworks, art historians typically favor the original form of the paintings and tend to view new technological solutions for processing them as potentially damaging, such as the circle transformation of the handscroll, leading to doubts and concerns. However, once our handscroll processing method is completed and presented, art historians realize the immense benefits of the new approach, evincing a high degree of interest and appreciation, and exhibiting readiness to embrace and adopt this technological solution. When engaging in interdisciplinary research, it is crucial to recognize that our new proposals may encounter resistance due to violating empirical habits. Therefore, it is important to comprehensively instruct and present the potential advantages of the new methods to collaborative researchers so that they can gain a thorough comprehension, ultimately appreciating the proposed solution.
Data credibility. Art historians place great importance on the reliability of data sources, as unreliable or unclear data sources empirically lead to unreliable or even wrong conclusions. This necessitates a substantial amount of effort to verify the data, diverting attention away from more in-depth research pursuits. In contrast, the proposed ScrollTimes system has undergone comprehensive verification and supplementation from multiple sources. Moreover, the system also includes detailed information on data sources, origins, and source database links. Such features can significantly bolster art historians’ trust in the data, rendering them more amenable to utilizing the system.
Data complementarity. Most of the databases and systems operate unidirectionally, with data providers solely responsible for managing and distributing data to researchers. However, our system features a bidirectional approach that enables researchers not only to access data from the database but also to contribute additional information through analysis to enhance the database. This approach facilitates mutual data complementation and reinforcement, resulting in a more robust and comprehensive database.
Expansion of research fields. Our work establishes a workflow in the field of handscrolls that leverages multiple models to process and integrate diverse data sources to gain profound insights. This approach can be expanded to research other domains of cultural heritage. For example, in the investigation of ancient architectural heritage, in addition to the information from the architecture itself, various multimodal sources such as related books, paintings, poems, etc. can also be used for verification and supplementation.
@@ 7.2 Limitations and Future Work
Data source expansion. Although our ScrollTimes system has already integrated numerous databases, we are eager to incorporate additional data sources, particularly those related to aesthetic documents, such as “Empty and Full: The Language of Chinese Painting” [15] and “The Chinese on the Art of Painting” [58] and so on. However, these data are merely recorded as raw texts, lacking a systematic and structured database format, which makes it challenging to use them for verification and supplementation.
Painting biography comparison. The ScrollTimes system does not provide an intuitive comparison of painting biographies, and art historians have suggested that comparing the biographies of different paintings can reveal cohort-level cultural heritage activities, thereby enabling analysis and comparisons of heritage styles. Those shortcomings highlight opportunities for improvement in future work.
@@ 8 Conclusion
To thoroughly study handscrolls, an art form that contains rich historical information, we present a visual analysis system called ScrollTimes, which tracks the origins and information contained in handscrolls by linking multiple data sources. We extract, verify, and supplement the elements and information contained in handscrolls, and create an interactive biography to further reveal their essence. Our approach offers a window into history, fostering a holistic understanding of handscroll provenance and historical significance. Two case studies and expert interviews demonstrate the usefulness and effectiveness of our system. We have summarized the lessons learned from developing ScrollTimes and believe that these implications will guide the system that is more comprehensive in functionality, richer in data sources, and more effective as a tool for art historians.

### v2

@@ 6.3 Art Historian Reviews
Effectiveness of the system. All experts agreed that ScrollTimes significantly benefits painting provenance research, particularly in seal identification and inscription analysis. A4 highlighted how it streamlines their workflow, transforming a “manual workshop” process into a productivity tool, freeing them to focus on more analytical tasks. A3 emphasized the time-saving aspect of integrating multiple databases and how ScrollTimes helps pinpoint research entry points. Additionally, A1, A2 and A8 expressed enthusiasm for ScrollTimes as a teaching aid in art history classrooms to enhance efficiency and student engagement.
Usability of the system. All the experts confirmed the usability of the system. A5 commented, “ScrollTimes allows us to visually access handscroll elements features, which was not previously feasible.” A3-5 appreciated the eye-catching circular design of the handscroll (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryA). While the circular format may deviate from the original handscroll’s aesthetic, A4 argued that it offers a holistic view of the entire handscroll, and the interaction can preserve the original information. A8 stated that the Linked Data View (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryB) provided valuable social network information. A2 and A3 mentioned that the Cohort View (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryB2) was unfamiliar to them and required some learning effort. A1 and A4 appreciated the Biography View (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryC), finding it a compelling metaphor representing life through a handscroll. They saw the encoded information as rich and potentially revealing new insights for analysis. A3 and A8 commented, “the Uncertainty Data View (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryE) opens up new research perspectives on less common historical figures for us.” A3, A5, and A6 appreciated the Similar Painting View (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryD), believing it could assist them in observing stylistic transmissions. A7 stated, “The exploration workflow makes sense to me; it aligns with the traditional workflow and offers numerous useful interactions.”
@@ 7 Discussion
This section summarizes lessons learned from working with art historians (Section 3.1) and discusses limitations and future work.
@@ 7.1 Design Implications
Open-mindedness. When art historians encounter a new design that surpasses their conventional understanding, it is necessary to showcase the advantage of the new design and preserve the traditional analytical methods of art historians. When engaging in interdisciplinary research, it is crucial to recognize that our new proposals may encounter resistance due to violating empirical habits. Therefore, it is important to comprehensively instruct and present the potential advantages of the new methods to collaborative researchers so that they can gain a thorough comprehension, ultimately appreciating the proposed solution.
Provenance as narratives. Historians typically specialize in specific periods for an in-depth understanding. Examining historical events of individuals [82] or cohorts [81] provides a glimpse into short-term historical viewpoints. Through ScrollTimes, we propose to interweave these fragmented viewpoints with cultural artifacts, much like intertwining colorful threads, thereby providing a broader perspective into the evolution of cultural understanding throughout history. Integrating micro (historical figures) and macro (paintings) perspectives offers new narratives on the long-lasting impact of the short-lived individuals.
Generalizability. Our workflow for handscrolls leverages multiple models to process and integrate diverse data sources to gain profound insights. Moreover, the seam-carving technique can accommodate various handscroll sizes and is intended to inspire more creative designs on the traditional art forms. Its capability in resizing paintings and retaining important details can be transferred to other artefacts to align with specific requirements. Other cultural heritage domains, such as the investigation of ancient architectural heritage, can be verified and supported with various multimodal sources (e.g., literature and poems).
@@ 7.2 Limitations and Future Work
Data source expansion. Although ScrollTimes has already integrated numerous databases, we are eager to incorporate additional data sources, particularly those related to aesthetic documents, such as “Empty and Full: The Language of Chinese Painting” [14] and “The Chinese on the Art of Painting” [69] and so on. However, these data are recorded as raw texts, lacking a structured database format, making verification and supplementation challenging.
Painting biography comparison. The ScrollTimes system does not provide an intuitive comparison of painting biographies. Comparing different biographies can reveal cohort-level cultural heritage activities, thereby enabling analysis and comparisons of heritage styles. This will, however, introduce more elements to be compared. An underexplored direction is to compare paintings in larger displays (e.g., immersive devices [40, 45] and combining multiple displays [72]).
@@ 8 Conclusion
This research explores the provenance of handscrolls, representative cultural artifacts with extensive historical content. We collaborated with art historians to develop ScrollTimes, which employs a three-tiered methodology (artifact, contextual, and provenance levels) for creating handscroll “Biography.” We extract, validate, and enhance elements within handscrolls by integrating data from multiple sources, creating an interactive biography to explore their historical significance. Two case studies and interviews have demonstrated the system’s utility. In the future, we aim to improve the system for simultaneous handscroll biography comparisons and expand its use to trace the provenance of various cultural artifacts using the three-tiered methodology.
Acknowledgements.
The authors wish to thank Peiyi Jiang, Hangchao Gong and Ollie Woodman at Zhejiang University, the editors at CBDB, and Thomas P. Kelly and Yukio Lippit at Harvard University for their valuable feedback and contribution to this project. This work was supported by the Fundamental Research Funds for the Central Universities (226-2022-00235), National Natural Science Foundation Key Project (62132017).